OpenStudy (anonymous):

Your urn contains 7 red balls, and 8 blue balls, and you draw 4 balls. In how many ways can the selection be made so that at least one of each color is drawn?

OpenStudy (dumbcow):

15C4 - 7C4 - 8C4 = 1365 - 35 - 70 = 1260